# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

### v3

@@ Introduction
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a mechanism for generalist behaviors.
@@
Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization. To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world navigation environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes. The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling.
@@
Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
@@
  • No context in training (Task 1) - a relation that the agent did not come across during training. It is simulated by placing the goal inside an empty box. During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but boxes contained keys.
  • Different context in testing (Task 2) - using a relation that the agent encountered during training but in a new context during testing. It is simulated by locking doors in the central room and placing keys directly in the room. During training, doors in the central room were never locked, nor was there direct access to keys.
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms. It is only done once, not repeatedly.
Learning to interact using type space knowledge
@@
Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
@@
Parameter Value Parameter Value
meta exp size 8 gae lambda 0.95
exp size 64 entropy coeff 1e-04
meta minibatch size 64 value coeff 1.0
minibatch size 64 meta entropy coeff 1e-04
meta learning rate 1e-05 meta value coeff 1.0
learning rate 1e-05 ppo clip 0.2
gamma 0.99 ppo steps 4
@@
Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate learning both the meta-level and interactions such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance space graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-order relations and apply these learned relations and interactions with objects to solve problems.
@@
Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level. Since the agent’s environmental return is determined by the number of steps taken to complete a task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of the value of meta-level interaction recommendations for agents. We evaluate the distribution of environmental returns for each agent across four tasks. For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked.
Figure 5: Evaluation return profiles of agents performing various tasks show that the proposed agents benefit from meta-level guided interactions.
To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$. Figure 5 shows the return profiles of agents. It is evident from the distribution of returns that the proposed agents achieve higher returns on all tasks compared to the base agent, including dynamic tasks like Task 3. This demonstrates that meta-guided interactions not only assist agents in completing tasks in fewer steps, but also enable them to perform a wider range of tasks in both static and dynamic scenarios. Furthermore, higher returns of the second variant show that agents benefit from treating reachability as a separate interaction. In the absence of separate reachability, reaching is embedded into interactions. Separate reachability allows shared learning followed by interactions, which is particularly beneficial in dynamically shifting scenarios.
@@
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a door that is obstructed by a ball and locked. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. During evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment itself. Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
@@ Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, type-space graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects. Second, rather than acting directly in the environment, interactions with objects are critical in performing tasks. To accomplish this, we introduced an intermediary interaction level. The agent’s meta-level interprets abstract relational knowledge, learns a meta-policy using type-space graphs, and makes recommendations on which objects to interact with and how to interact with them. In this sense, our model functions as an interpretation-based interaction goal-generating machine. Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries. We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization. Lastly, Our framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels. This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.

### v4

@@ Introduction
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a metacognitive mechanism for knowledge learning, abstract reasoning and generalist behaviors.
@@
We aim to examine whether agents can learn to associate abstract relational knowledge with object and interaction types, leverage this knowledge for interaction learning and guide interactions to exhibit flexible behavior when solving complex tasks. To assess the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world navigation environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes. The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling.
@@
Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s knowledge learning and flexible task solving capabilities, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
@@
  • No context in training (Task 1) – a relation that the agent did not come across during training. It is simulated by placing the goal inside an empty box. During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but the boxes contained keys.
  • Different context in testing (Task 2) – using a relation that the agent encountered during training but in a new context during testing. It is simulated by locking doors in the central room and placing keys directly in the room. During training, the doors in the central room were never locked, nor was there direct access to the keys.
  • Dynamic changes in the environment (Task 3) – when some event occurs, the environment changes. During training, the goal location in an episode remained unchanged. To simulate this, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms. It is done only once, not repeatedly.
Knowledge Learning to Interact
@@
Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes such as color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
@@
Parameter Value Parameter Value Parameter Value
meta exp size 8 learning rate 1e-05 gamma 0.99
exp size 64 entropy coeff 1e-04 gae lambda 0.95
meta minibatch size 64 value coeff 1.0 ppo clip 0.2
minibatch size 64 meta entropy coeff 1e-04 ppo steps 4
meta learning rate 1e-05 meta value coeff 1.0
@@
Figure 4: Agents learning to interact : The learning progress of both the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, demonstrating the learning of meta-level as well as interaction policies such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-order relations and apply these learned relations in interactions with objects to solve problems.
@@
Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level. Since the agent’s environmental return is determined by the number of steps taken to complete a task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of the value of meta-level interaction recommendations for agents. We evaluate the distribution of environmental returns for each agent across four tasks.
Figure 5: Evaluation return profiles of agents performing tasks show that the agents benefit from meta-level guided interactions.
For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked. To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$. Figure 5 shows the return profiles of agents. It is evident from the distribution of returns that the proposed agents achieve higher returns on all tasks compared to the base agent, including dynamic tasks like Task 3. This demonstrates that meta-guided interactions not only assist agents in completing tasks in fewer steps, but also enable them to perform a wider range of tasks in both static and dynamic scenarios. Furthermore, higher returns of the second variant show that agents benefit from treating reachability as a separate interaction. In the absence of separate reachability, reaching is embedded into interactions. Separate reachability allows shared learning followed by interactions, which is particularly beneficial in dynamically shifting scenarios.
@@
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a locked door that is obstructed by a ball. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. In evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment. Figure 6 shows evaluation return profiles for the base task, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
@@ Conclusion
Artificial intelligence systems must develop knowledge-centric metacognitive abilities in the pursuit of abstract reasoning, knowledge learning, adaptive behavior and autonomy. The proposed metacognitive approach promotes these capabilities through three core contributions. First, knowledge learning and adaptation necessitate abstraction and composition while preserving relational knowledge structures. In our model, type graphs are used to represent and ground natural abstractions while preserving relational knowledge among objects. Second, rather than acting directly in the environment, interactions with objects are essential in task execution. To accomplish this, we introduced an intermediary interaction level. At the meta level, the agent interprets abstract relational knowledge, learns a meta-policy over type graphs, and generates recommendations regarding which objects to interact with and how. In this sense, the proposed model functions as an interpretation-based interaction goal generating machine. Third, meta level guided interactions executed through interaction policies provide the loose coupling necessary for generalization by acting as intermediaries. The metacognitive approach facilitates knowledge learning of policies, knowledge guided interactions aid generalization and transferable interaction policies. Lastly, the proposed framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels. It is particularly promising for physical-world domains where abstract relational knowledge and compositional interactions can be exploited, such as robotics, autonomous, and cognitive systems.